# World-Env: Leveraging World Model as a Virtual Environment for VLA Post-Training

### v3

@@ 4.2 VLM-guided Instant Reflector
Previous methods (Tan et al., 2025; Lu et al., 2025) rely on simulators to provide binary task success signals, using sparse discrete rewards for RL post-training. These approaches suffer from a key limitation: the lack of termination-aware feedback, causing policies to often continue executing redundant actions after task completion (e.g., over-scooping after object placement). To address this, we propose a VLM-guided instant reflector that leverages LLaVA (Liu et al., 2023c), a pretrained vision-language model, to provide a continuous-valued reward signal.
@@
For training, we utilize per-frame binary success labels: for each trajectory, every timestep $t$ is annotated with $y_{t}\in\{0,1\}$, indicating whether the task is completed at or before $t$. These labels are derived from two sources: (1) expert trajectories from the LIBERO dataset (Liu et al., 2023a), where success is determined by task-specific criteria, and (2) policy-generated trajectories collected in simulator (Section 4.1), labeled using an oracle that monitors ground-truth task states in the simulator. The reward head $\mathcal{R}_{\theta}$ is trained with binary cross-entropy (BCE) loss:
@@
This supervision enables the reflector to recognize task completion as soon as it occurs, rather than relying on trajectory-level signals. During RL, we use the reward sparsely: the return is computed using a single reward assigned at the termination timestep (or at $T$ if no termination occurs), ensuring compatibility with standard policy gradient methods. This design allows World-Env to simultaneously support real-time termination and efficient policy learning, effectively addressing the execution inefficiency of prior VLA post-training approaches (Tan et al., 2025; Lu et al., 2025). More details can be found in the supplementary material.
@@ 4.3 Post Training of VLA Model
Our reinforcement learning pipeline employs a PPO-style objective with continuous reward signals; the full algorithm is provided in the supplementary material. Following Tan et al. (2025), training proceeds in three stages: rollout generation, advantage estimation, and policy optimization.
During rollout, we generate trajectories $\tau=(\mathbf{o}_{1:T},\mathbf{s}_{1:T},\mathbf{g},\mathbf{a}_{1:T})$ using the world simulator (Section 4.1). Starting from an initial observation $\mathbf{o}_{1}$, proprioceptive state $\mathbf{s}_{1}$, and language instruction $\mathbf{g}$, the deterministic VLA policy $\pi_{\theta}$ predicts a base action $\bm{\mu}_{t}=\pi_{\theta}(\mathbf{o}_{1:t},\mathbf{s}_{1:t},\mathbf{g})$. A separate scale head, trained to model action uncertainty, outputs a log-scale parameter $\bm{\beta}_{t}$. Together, they define a factorized Laplace distribution, from which the executed action $\mathbf{a}_{t}$ is sampled. This enables adaptive, uncertainty-aware exploration. The world simulator then predicts the next observation $\mathbf{o}_{t+1}$ using proprioceptive state $\mathbf{s}_{t+1}$. The VLM-guided instant reflector evaluates the partial visual trajectory $\mathbf{o}_{1:t+1}$ and outputs a step-wise reward $R(\mathbf{o}_{1:t+1},\mathbf{g})\in[0,1]$. Rollout terminates either at the maximum timestep $T$ or when $R(\mathbf{o}_{1:t+1},\mathbf{g})>\eta$. For RL, we assign a single trajectory-wise reward $R_{n}=R(\mathbf{o}_{1:t_{\text{end}}},\mathbf{g})$, where $t_{\text{end}}$ is the termination or final timestep.
We adopt Leave-One-Out Proximal Policy Optimization (LOOP) (Chen et al., 2025) that combines RLOO (Ahmadian et al., 2024) based advantage estimation and PPO (Schulman et al., 2017) for policy updating. For each initial state, we generate $N$ rollouts $\{\tau_{1},\dots,\tau_{N}\}$ using a fixed behavior policy $\pi_{\phi}$ (the policy at the beginning iteration). Each trajectory receives a scalar reward $R_{n}$ from the instant reflector. The RLOO baseline for trajectory $n$ is the average reward of the other rollouts:
@@
$\mathcal{L}_{\text{PPO}}=-\frac{1}{\sum_{n}T_{n}}\sum_{n=1}^{N}\sum_{t=1}^{T_{n}}\min\left(r_{t,n}A_{n},\,\text{clip}(r_{t,n},1-\epsilon,1+\epsilon)A_{n}\right),$ (5)
with $T_{n}$ denotes the length of trajectory $n$ and $\epsilon$ refers to the clipping threshold. Note that the advantage $A_{n}$ is broadcasted to all timesteps within trajectory.
Unlike prior methods that use binary rewards ($R\in\{0,1\}$) and require balanced success/failure rollouts for stable training, our continuous reward signal ($R\in[0,1]$) provides finer-grained feedback. This helps enhance rollout efficiency and training stability, particularly in data-scarce settings.
@@ 4.4 Implementation Details
All our experiments are conducted on 8 NVIDIA H20 GPUs (96 GB memory each). We adopt LoRA (Hu et al., 2022) with rank 32 for parameter-efficient fine-tuning of the vision-language backbone, while the action head and scale head are trained with full parameters. We use a batch size of 4. The LoRA adapters are optimized with a learning rate of $1\times 10^{-4}$, and the action/scale heads are trained with a learning rate of $1\times 10^{-5}$. We set the number of rollouts per iteration to $N=8$ and the PPO clipping threshold to $\epsilon=0.1$.
$\pi_{0}$ (Black et al., 2024) 67.6 68.4 80.2 28.2 61.1
$\pi_{0}$+FAST (Pertsch et al., 2025) 59.2 76.8 59.2 24.8 55.0
OpenVLA (Kim et al., 2024) 73.2 55.0 82.4 32.2 60.7
UniVLA (Bu et al., 2025) 82.0 76.2 84.4 56.4 74.75
OpenVLA-OFT (Kim et al., 2025) 84.0 74.2 84.2 57.0 74.85
OpenVLA-OFT + Post-training (Ours) 86.4 86.6 87.6 57.8 79.6
@@ 5 Experiments
Figure 3: Comparison between our method and SFT on multi-goal tasks. Note, all results are collected every 5 training steps for three distinct goals.
Extra Data Reward Head LIBERO Goal LIBERO Object LIBERO Spatial LIBERO Long
68.4 75.2 73.2 42.2
79.8 81.8 78.4 44.6
68.8 76.4 74.4 43.8
86.4 86.6 87.6 57.8
Benchmark.
We evaluate our model on the LIBERO benchmark (Liu et al., 2023a), a simulation-based robotic learning platform designed for vision-language manipulation tasks. The benchmark includes four task suites targeting distinct cognitive challenges: LIBERO-Spatial focusing on spatial reasoning via object arrangement; LIBERO-Goal assessing goal-conditioned planning with end-state requirements; LIBERO-Object testing object-centric manipulation across categories; LIBERO-10 (LIBERO-Long) addressing prolonged sequential decision-making. Each suite contains 10 tasks with 50 trajectories for training and 50 for testing per task; we train OpenVLA-OFT using only 5 trajectories from the training split to validate performance under extreme data scarcity, while evaluating on the full trajectory test split to demonstrate the generalization capability.
Baselines.
We compare our method with five state-of-the-art VLA frameworks including $\pi_{0}$ (Black et al., 2024), $\pi_{0}$ + FAST (Pertsch et al., 2025), OpenVLA (Kim et al., 2024), UniVLA (Bu et al., 2025), and OpenVLA-OFT (Kim et al., 2025). All methods are trained with standard supervised fine-tuning (SFT). For fair evaluation, all baselines are retrained under identical 5-trajectory per-task constraints, with performance metrics reported on the complete test set.
@@ 5.1 Comparison with State-of-the-art methods
Table 1 presents success rate comparison between our method and the baseline models. As shown, our method gains higher task success rate, demonstrating the effectiveness of our proposed post-training strategy. Figure 3 further compares our method and the supervised fine-tuning (SFT) baseline on multi-goal tasks, where we can see that our approach achieves superior performance within only 20 training steps, clearly outperforming the compared SFT model. This rapid convergence and early dominance highlight the efficiency and effectiveness of our method in learning conditioned policies with minimal training iterations.
Figure 5: Real-world rendering results of world simulator. We show a video sequence generated by our world simulator in real-world scene.
5.2 Ablation studies
Effect of World Simulator.
We investigate how the generative capabilities of world simulator affect the performance of our method. Figure 4 evaluates two world simulator variants: (1) w/o extra: trained solely on human-annotated successful trajectories, (2) w/ extra: enhanced with our collected data containing both successful and failed trajectories. As shown, the model trained without extra data struggles with object tracking, particularly when the VLA model’s action predictions deviate from ideal trajectories. This is because the world simulator only observes successful interactions during training, making it unable to simulate complex object states caused by suboptimal actions. In contrast, our model demonstrates significant improvements in robotic arm tracking precision and object interaction fidelity. Table 2 quantitatively validates these observations: when the world simulator generates low-quality images, VLA training effectiveness drops. This correlation highlights the importance of diverse training data in building robust world simulator that can handle real-world action variations. Figure 5 further manifests the simulator’s ability to generate photorealistic observations with accurate physical interactions, showing that our method can be adopted in real-world scene application. Please see also the supplementary material for video results.
Effect of Instant Reflector.
We investigate the effect of instant reflector on our framework’s performance. As summarized in Table 2, we perform evaluation for two strategies: (1) w/o reward head: Direct use of pre-trained VLMs with prompt-based binary classification (yes/no) (2) w/ reward head: We integrate a trainable reward head that scores action sequences on a continuous scale. The difference lies in how each approach assesses alignment between generated video sequences and text instructions. While the baseline leverages VLMs’ inherent language understanding capabilities through fixed prompts, our method explicitly trains the reward head to quantify task completion progress via scalar scores. Experimental validation demonstrates that a naive use of pre-trained VLMs brings weak performance gains and may degrade VLA model learning in complex scenarios. This limitation stems from the mismatch between VLMs’ general language vision alignment and the specific action evaluation requirements. In contrast, our reward head, trained on diverse success/failure trajectories, achieves higher accuracy in distinguishing successful actions.
Effect of Termination Signals.
Table 3 further validates the effectiveness of our task success detection capability. While conventional methods rely on simulator-provided termination signals due to their inability to assess task completion, our approach employs a VLM-guided instant reflector that dynamically evaluates task success and enables early termination upon achievement. To verify this advantage, we set all compared baseline methods to strictly follow the maximum step limit for termination, whereas our framework utilizes instant reflector predictions as stopping criteria. As shown in Table 3, the compared baseline methods exhibit clear performance degradation under this setting because redundant post-success actions from delayed termination may disrupt object states after task completion (see Figure 6). In contrast, our method avoids such interference by terminating execution immediately upon detecting success signals, demonstrating our instant reflector’s capacity to preserve task outcomes through timely stopping decisions.
@@ 5.2 Ablation studies
Figure 6: Post-success failure in VLA execution. An illustrative example for “put the wine bottle on top of the cabinet” shows the VLA model completes the task (frames a-b), but fails due to delayed termination (frames c-e), validating the necessity of dynamic termination mechanism.
@@
$\pi_{0}$ (Black et al., 2024) 55.4 71.0 72.6 20.6 54.9
$\pi_{0}$+FAST (Pertsch et al., 2025) 21.2 74.0 44.8 15.0 38.75
OpenVLA (Kim et al., 2024) 68.4 47.4 59.8 26.6 50.55
UniVLA (Bu et al., 2025) 72.0 75.2 66.4 48.0 65.4
OpenVLA-OFT (Kim et al., 2025) 67.4 73.8 71.2 39.8 63.05
@@ 5.3 Limitations and future work
Despite the effectiveness of our method in enhancing VLA manipulation capabilities, it still has the following limitations. First, the performance of both world simulator and instant reflector depends on massive training data to achieve high-fidelity simulation and accurate task evaluation. Second, our VLA model optimization is slower than concurrent methods due to world simulator generation bottlenecks. In the future, we will focus on addressing these limitations.
@@ 6 Conclusion
We present a post-training framework World-Env for Vision-Language-Action (VLA) models that eliminates reliance on physical environment interaction. We introduce three core innovations: (1) RL post-training by exploration in World-Env enables policy refinement, achieving strong performance with only 5 demonstrations per task; (2) exploration in World-Env reduces physical experimentation costs; and (3) dynamic termination via VLM-guided instant reflector prevents redundant post-success actions. Experimental validation on complex manipulation tasks demonstrates our method’s superiority in low-data regimes.
Appendix A Algorithm
@@
10:      Predict base action: $\bm{\mu}_{t}\leftarrow\pi_{\phi}(\mathbf{o}_{1:t},\mathbf{s}_{1:t},\mathbf{g})$
11:      Predict log-scale: $\bm{\beta}_{t}\leftarrow\beta_{\phi}(\mathbf{o}_{1:t},\mathbf{s}_{1:t},\mathbf{g})$
12:      Sample action: $\mathbf{a}_{t}\sim\text{Laplace}(\bm{\mu}_{t},\exp(\bm{\beta}_{t}))$
@@
31:   Update $\pi_{\theta}$ and $\beta_{\theta}$ by minimizing PPO loss (Eq. 5)
@@
Appendix B More Implementation Details
B.1 Deatils of Scale Head
Our method builds upon OpenVLA-OFT (Kim et al., 2025), which predicts continuous actions via an action head that takes hidden states $f\in\mathbb{R}^{d}$ as input and employs L1 loss for action regression:
$\mathcal{L}_{\text{L1}}=\|\mathbf{a}_{\text{gt}}-\bm{\mu}\|_{1}\quad\text{where }\bm{\mu}=\text{MLP}_{\text{action}}(f).$ (6)
To model heteroscedastic uncertainty in action prediction, we introduce a scale head with the same MLP architecture as the action head, as shown in Figure 7. This scale head outputs log-scale parameters $\bm{\beta}$ through:
$\bm{\beta}=\text{MLP}_{\text{scale}}(h),$ (7)

### v5

@@ 4.2 VLM-Guided Instant Reflector
Previous methods [44, 36] rely on simulators to provide binary task success signals, using sparse discrete rewards for RL post-training. These approaches suffer from a key limitation: the lack of termination-aware feedback, causing policies to often continue executing redundant actions after task completion (e.g., over-scooping after object placement). To address this, we propose a VLM-guided instant reflector that leverages a pretrained vision-language model to provide a continuous-valued reward signal.
@@
For training, we utilize per-frame binary success labels: for each trajectory, every timestep $t$ is annotated with $y_{t}\in\{0,1\}$, indicating whether the task is completed at or before $t$. These labels are derived from two sources: (1) expert trajectories from the LIBERO dataset [33], where success is determined by task-specific criteria, and (2) policy-generated trajectories collected in simulator (Section 4.1). The reward head $\mathcal{R}_{\theta}$ is trained with BCE loss:
@@
This supervision enables the reflector to recognize task completion as soon as it occurs. During RL, we use the reward sparsely: the return is computed using a single reward assigned at the termination timestep (or at $T$ if no termination occurs).
This design enables World-Env to support real-time termination while achieving both high sample efficiency and stable policy learning. In contrast, prior VLA post-training approaches [44, 36] rely on binary rewards (e.g., 1 for success, 0 for failure), which lead to degenerate advantage estimates when rollout trajectories are homogeneous, i.e., when all episodes succeed or all fail. In such cases, the empirical advantage values collapse to zero across the batch, providing no learning signal for policy updates and drastically reducing training efficiency. By instead employing a continuous reward signal in the range $[0,1]$ that reflects fine-grained task progress, our method ensures non-trivial advantage estimates. This eliminates the need to balance successful and failed rollouts during data collection, enabling more efficient use of samples.
@@ 4.3 Post Training of VLA Model
Our reinforcement learning pipeline employs a RLOO [1] objective with continuous reward signals; the full algorithm is provided in the supplementary material. Following [44], training proceeds in three stages: rollout generation, advantage estimation, and policy optimization.
During rollout, we generate trajectories $\tau=(\mathbf{o}_{1:T},\mathbf{s}_{1:T},\mathbf{g},\mathbf{a}_{1:T})$ using the world simulator (Section 4.1). Starting from an initial observation $\mathbf{o}_{1}$, proprioceptive state $\mathbf{s}_{1}$, and language instruction $\mathbf{g}$, the deterministic VLA policy $\pi_{\theta}$ predicts a base action $\boldsymbol{\mu}_{t}=\pi_{\theta}(\mathbf{o}_{1:t},\mathbf{s}_{1:t},\mathbf{g})$. A separate scale head, trained to model action uncertainty, outputs a log-scale parameter $\boldsymbol{\beta}_{t}$. Together, they define a factorized Laplace distribution, from which the executed action $\mathbf{a}_{t}$ is sampled. This enables adaptive, uncertainty-aware exploration. The world simulator then predicts the next observation $\mathbf{o}_{t+1}$ using proprioceptive state $\mathbf{s}_{t+1}$. The VLM-guided instant reflector evaluates the partial visual trajectory $\mathbf{o}_{1:t+1}$ and outputs a step-wise reward $R(\mathbf{o}_{1:t+1},\mathbf{g})\in[0,1]$. Rollout terminates either at the maximum timestep $T$ or when $R(\mathbf{o}_{1:t+1},\mathbf{g})>\eta$. For RL, we assign a single trajectory-wise reward $R_{n}=R(\mathbf{o}_{1:t_{\text{end}}},\mathbf{g})$, where $t_{\text{end}}$ is the termination or final timestep.
We adopt Leave-One-Out Proximal Policy Optimization (LOOP) [11] that combines RLOO [1] based advantage estimation and PPO [43] for policy updating. For each initial state, we generate $N$ rollouts $\{\tau_{1},\dots,\tau_{N}\}$ using a fixed behavior policy $\pi_{\phi}$ (the policy at the beginning iteration). Each trajectory receives a scalar reward $R_{n}$ from the instant reflector. The RLOO baseline for trajectory $n$ is the average reward of the other rollouts:
@@
$\mathcal{L}_{\text{PPO}}=-\min(r_{t,n}A_{n},\,\text{clip}(r_{t,n},1-\epsilon,1+\epsilon)A_{n}),$ (6)
with $\epsilon$ representing to the clipping threshold. Note that the advantage $A_{n}$ is broadcasted to all timesteps.
Table 2: Comparison with simulator-based RL method.
Method Goal Object Spatial Long
RIPT-VLA [44] 86.2 83.4 88.6 58.4
Ours 86.4 86.6 87.6 57.8
Figure 7: Real-world experiment on task “clean table”. Full video result can be found in the supplementary material.
@@ 4.4 Implementation Details
Experiments were performed on 8 NVIDIA H20 GPUs (96 GB VRAM each), with a total training time of approximately 48 hours. We adopt LoRA [22] with rank 32 for parameter-efficient fine-tuning of the vision-language backbone, while the action head and scale head are trained with full parameters. We use a batch size of 4. The LoRA adapters are optimized with a learning rate of $1\times 10^{-4}$, and the action/scale heads are trained with a learning rate of $1\times 10^{-5}$. We set the number of rollouts per iteration to $N=8$ and the PPO clipping threshold to $\epsilon=0.1$.
Method Clean table Put green toy Put red toy Put orange toy
OpenVLA-OFT 20 30 30 20
Ours 30 50 40 50
Table 3: Comparison results of real-world experiments.
@@ 5 Experiments
Benchmarks. We evaluate our model on the LIBERO benchmark [33], a simulation-based robotic learning platform designed for vision-language manipulation tasks. The benchmark includes four task suites targeting distinct cognitive challenges: LIBERO-Spatial focusing on spatial reasoning via object arrangement; LIBERO-Goal assessing goal-conditioned planning with end-state requirements; LIBERO-Object testing object-centric manipulation across categories; LIBERO-10 (LIBERO-Long) addressing prolonged sequential decision-making. Each suite contains 10 tasks with 50 trajectories for training and 50 for testing per task; we train OpenVLA-OFT using only 5 trajectories from the training split to validate performance under extreme data scarcity, while evaluating on the full trajectory test split to demonstrate the generalization capability.
Baselines. We compare our method with five state-of-the-art supervised fine-tuning (SFT) methods including $\pi_{0}$ [6], $\pi_{0}$ + FAST [39], OpenVLA [27], UniVLA [9], OpenVLA-OFT [26] and one simulator-based RL method RIPT-VLA [44]. For fair evaluation, all baselines are retrained under identical 5-trajectory per-task constraints, with performance metrics reported on the complete test set.
@@ 5.1 Comparison with State-of-the-art Methods
Table 1 presents success rate comparison between our method and the baseline models. As shown, our method gains higher task success rate, demonstrating the effectiveness of our proposed post-training strategy. Figure 5 further compares our method and the supervised fine-tuning (SFT) baseline on multi-goal tasks, where we can see that our approach achieves superior performance within only 20 training steps, clearly outperforming the compared SFT model. This rapid convergence and early dominance highlight the efficiency and effectiveness of our method in learning conditioned policies with minimal training iterations. Table 2 presents a comparison with RIPT-VLA [44], a simulator-based reinforcement learning approach. Our method achieves comparable performance while offering a crucial practical advantage: it is readily deployable in real-world settings. In contrast, RIPT-VLA is limited to simulated environments.
@@
Figure 8: Post-success failure in VLA execution. An illustrative example for “put the wine bottle on top of the cabinet” shows the VLA model completes the task (a-b), but fails due to delayed termination (c-e), validating the necessity of dynamic termination mechanism.
@@
$\pi_{0}$ [6] 55.4 71.0 72.6 20.6 54.9
$\pi_{0}$+FAST [39] 21.2 74.0 44.8 15.0 38.75
OpenVLA [27] 68.4 47.4 59.8 26.6 50.55
UniVLA [9] 72.0 75.2 66.4 48.0 65.4
OpenVLA-OFT [26] 67.4 73.8 71.2 39.8 63.05
@@
Extra Data Reward Head Goal Object Spatial Long
68.4 75.2 73.2 42.2
79.8 81.8 78.4 44.6
68.8 76.4 74.4 43.8
86.4 86.6 87.6 57.8
5.2 Real-world Experiment
We conduct real-world experiments across four tasks, i.e., “clean table”, “put green toy in cabinet”, “put red toy in cabinet”, and “put orange toy in cabinet”, Figure 7 shows several sampled frames from the “clean table” task, where the goal is to pick up three toys and place them into the bucket. For each task, we collect 10 trajectories for both policy SFT and world model fine-tuning. As shown in the table 3, our method outperforms OpenVLA-OFT on all four tasks, validating that our framework can be effectively transferred to real-world environments.
5.3 Ablation Studies
Effect of World Simulator. We study how the world simulator’s generative quality affects performance. Figure 6 compares two variants: (1) w/o extra: trained only on successful human trajectories; and (2) w/ extra: augmented with our data containing both success and failure cases. Without exposure to suboptimal actions, the former fails to model resulting object states, leading to poor tracking under VLA prediction errors. Our enhanced simulator substantially improves arm tracking and interaction fidelity. Table 5 confirms that low-fidelity simulation degrades VLA training. Additional analysis, ablations, and videos are in the supplementary material.
Effect of Instant Reflector. Table 5 compares two variants: (1) w/o reward head: using a pre-trained VLM with prompt-based binary classification; and (2) w/ reward head: a trainable module scoring actions on a continuous scale. Experiments show that off-the-shelf VLMs offer limited gains and can impair learning in complex tasks due to misalignment with fine-grained action evaluation. In contrast, our reward head provides more accurate and reliable action assessment.
Effect of Termination Signals. Table 4 further validates our method’s ability to detect task success. Unlike conventional approaches, which rely on provided termination signals, our framework uses a VLM-guided instant reflector to dynamically assess completion and stop execution promptly. To isolate this advantage, we force all baselines to adhere strictly to the full horizon, while our method terminates based on reflector predictions. As shown, baselines suffer performance degradation due to redundant post-success actions that perturb object states (see Figure 8). In contrast, our approach preserves task outcomes by halting immediately upon success detection, demonstrating the reflector’s effectiveness in enabling timely termination.
@@ 5.4 Limitations and Future Work
Despite the effectiveness of our approach in enhancing VLA manipulation capabilities, several limitations remain. First, both the world simulator and the instant reflector rely on diverse training data to achieve high-fidelity simulation and accurate task evaluation. We anticipate that future advances in general-purpose world models will alleviate this dependency. Second, policy optimization in our framework is currently slower than in concurrent methods due to computational bottlenecks in simulator-based trajectory generation. Addressing these challenges through more efficient simulation will be a key focus of our future work.
@@ 6 Conclusion
We present World-Env, a post-training framework for Vision-Language-Action (VLA) models that eliminates the need for costly and potentially unsafe interactions with physical environments. By leveraging a physically consistent world simulator, enhanced with a geometry-aware feature injection strategy, World-Env enables safe and low-cost exploration in virtual settings. Furthermore, our VLM-guided instant reflector supports dynamic task-aware termination, preventing redundant post-success actions that could otherwise disrupt task outcomes. Extensive experiments on complex robotic manipulation tasks demonstrate that World-Env achieves superior performance, particularly in low-data regimes.
Acknowledgement. This work was supported by the National Natural Science Foundation of China (62471499), the Guangdong Basic and Applied Basic Research Foundation (2023A1515030002).
Supplementary Material
7 Algorithm
@@
10:          Predict base action: $\boldsymbol{\mu}_{t}\leftarrow\pi_{\phi}(\mathbf{o}_{1:t},\mathbf{s}_{1:t},\mathbf{g})$
11:          Predict log-scale: $\boldsymbol{\beta}_{t}\leftarrow\beta_{\phi}(\mathbf{o}_{1:t},\mathbf{s}_{1:t},\mathbf{g})$
12:          Sample action: $\mathbf{a}_{t}\sim\text{Laplace}(\boldsymbol{\mu}_{t},\exp(\boldsymbol{\beta}_{t}))$
@@
31:     Update $\pi_{\theta}$ and $\beta_{\theta}$ by minimizing PPO loss
@@
We show the full reinforcement learning post-training algorithm in Algorithm 1.
8 Comparison to Octo.
We compare with Octo [45] in Table 6. As shown in the table, our method outperforms Octo across all four LIBERO suites despite using only $10\%$ training data, manifesting its effectiveness.
Method LIBERO-Goal LIBERO-Object LIBERO-Spatial LIBERO-Long
Octo 84.6 85.7 78.9 51.1
Ours 86.4 86.6 87.6 57.8
Table 6: Quantitative comparison with Octo on LIBERO.
9 More Implementation Details
9.1 Deatils of Scale Head
Our method builds upon OpenVLA-OFT [26], which predicts continuous actions via an action head that takes hidden states $f\in\mathbb{R}^{d}$ as input and employs L1 loss for action regression:
$\mathcal{L}_{\text{L1}}=\|\mathbf{a}_{\text{gt}}-\boldsymbol{\mu}\|_{1}\qquad\text{where }\boldsymbol{\mu}=\text{MLP}_{\text{action}}(f).$ (7)
To model heteroscedastic uncertainty in action prediction, we introduce a scale head with the same MLP architecture as the action head, as shown in Figure 9. This scale head outputs log-scale parameters $\boldsymbol{\beta}$ through:
$\boldsymbol{\beta}=\text{MLP}_{\text{scale}}(h),$ (8)